Two cars traveling in opposite directions pass each other at 1:00 p.m. One of the cars travels at a speed of 90 km/h, while the other travels at a speed of 72 km/h. How far apart will they be at 2:00 p.m.?
A) 162 km B) 108 km C) 144 km D) 180 km
1
Expert's answer
2016-06-29T09:37:02-0400

The relative speed of the cars:

V = V1 + V2 = 90 + 72 = 162 km/h

By 2:00 pm they will be travelling for one hour. Therefore, the distance will be:

D = V x t = 162 x 1 = 162 km

Correct answer option: A) 162 km